OpenStudy (anonymous):

Sarah, Amy and Liam stand on some weighing scales two at a time. Here are the measurements: Sarah and Amy Sarah and Liam Liam and Amy 70kg 80kg 80kg How much does Liam weigh? show working please :)

OpenStudy (anonymous):

helpppppppp

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I don't know! sorry :( I love your new pic though :D

OpenStudy (anonymous):

lol thnaks re-re

OpenStudy (anonymous):

next time yhh :)

OpenStudy (mathmagician):

S+A=70 S+L=80 L+A=80 You see, that S+L=L+A => S=A S+A=70=>A=70/2=35 and finally L=80-35=45

OpenStudy (anonymous):

THNX MATH MAGICIAN UR A LIFE SAVER